wiki:UC20170814

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 8/14/17

Attending: Nancy Chanover, Mark Klaene, Rachel Kuzio de Naray, Mary Beth Kaiser, John Tobin, John Bally, Jon Holtzman, Mike Skrutskie, Chip Kobulnicky


Our new minute-taker, Anne Verbiscer, was unable to make this call. Thus, Nancy took the minutes for August.


User feedback and comments from institutional representatives:

  • Wyoming - Chip had nothing to report.
  • Washington - Scott was absent, but by email he had nothing new to report.
  • Virginia - Mike S. had nothing new to report.
  • Seattle - absent; no report
  • Oklahoma - John T. had nothing new to report.
  • NMSU - Jon H. had nothing new to report other than user distress about the unavailability of ARCES
  • NAPG - absent; no report
  • JHU - Mary Beth had nothing new to report.
  • Georgia State - Rachel reported that several of her users have been asking about the status of the new spectrograph. Nancy said that there has been some work going on behind the scenes concerning instrument design options and hiring an optical engineer. This instrument remains our highest priority initiative, and we hope to have a more detailed review schedule and build schedule by the ARC Board of Governors meeting in early November.
  • Colorado - John B. had nothing new to report.

Telescope and Instruments Report:

This call was held just as we were coming out of the 3.5m summer shutdown, so there was no formal telescope and instrument report this month. Instead, Mark gave a verbal update on the work that had been accomplished during the shutdown. A summary of that report is given below.

3.5-m Telescope and Instruments Highlights, 7/6/17 through 8/14/17

1) Telescope

It has been a productive albeit a somewhat frustrating summer shutdown. The primary mirror (M1) was washed. The tertiary mirror has a fresh coat of aluminum on it, but Mark and Bill are not happy with it even after having made 3 attempts at the realuminization. [For comparison, they did the ARCSAT mirror 6 times with similar results.] The coatings all have a slight fog pattern on them. This is something that happened to the tertiary a few years ago, which is why we are now aluminizing it every year. Bill and Mark did lots of work on the chamber and refined their procedures for cleaning the optics, pumping down the chamber, and firing, all to no avail; they had as good a pump down and firing as they could get but are still getting the fogging pattern. After several tests they are now convinced that something is etched in or on the glass that is not coming off with regular cleaning routines. They realuminized a few other small mirrors and they came out OK, which suggests that the problem is with the glass itself. They talked with people at MMT and other facilities and they are running out of local ideas and those provided by other observatories. For now the tertiary mirror is going back into the telescope in much better condition than when it was removed; it is definitely still improved although we are still not totally happy with the outcome and will continue to explore this issue with the ARCSAT mirror. The secondary mirror was not serviced; we will only do a recoating once we're happy with the results from the chamber. The actuators on the tertiary were rebuilt (all of the spares have been used up), and the telescope is going back together as we speak. There is not much collimation and on-sky engineering that will be required, so we expect things to go smoothly for the remainder of the week. Other preventive maintenance tasks were performed on the telescope; there were no problems or concerns.

2) Instruments

DIS will be going down and coming off the telescope in early September (9/4). Sarah Tuttle is coming out; she and Bill will open up DIS and replace O-rings and examine the contamination problem. The problem with the DIS cryotigers was a vacuum problem; the ion pumps were replaced and now it is functioning normally.

There was a minor hiccup with NICFPS but now it is functioning normally.

TripleSpec is back up and operational, thanks to the hard work of the UVA instrument team and local staff.

The ARCES problem with image readouts continues. Ed Leon and Ben Harris worked with Connie Rockosi (remotely) and safely removed the saddlebags from instrument and brought the instrument down to the lab where they could work on it more efficiently. They have seen the readout problem come and go; they believe that it is due to a cold solder joint or a connector that was susceptible. Several joints and connectors were replaced but the problem still comes and goes; they will resume troubleshooting later this week (among other summer shutdown tasks). Connie is tentatively set to come out the week of 8/21 in the event that ARCES won’t be fixed by end of this week. We are optimistic that the echelle will be back in service by the start of Q4, maybe even by September, but at this point we're not saying that it’s fixed. The instrument has been running under vacuum since they took the saddlebags off, so hopefully this will eliminate whatever contamination issues may have been lingering in the instrument.

Russet has continued her heroic efforts in modifying the schedule to account for the echelle downtime. She requested that the Users Committee reps help her track down echelle user candidates who might be able to use TripleSpec instead; that would be a natural trade since both instruments are normally scheduled during bright time.


Q3 3.5-m scheduling:

DIS users should be aware of the scheduled downtime in September (it is already built into the schedule).


NA2 Baffle Project

The observatory is exploring an initiative to redo the NA2 baffle, which would enable users to observe with the full 7.5' x 7.5' FOV of ARCTIC with no vignetting. Nancy will send out more information on this within the next several weeks.


TDAWG Update

The first TDAWG telecon is scheduled for August 28 at 11 am PDT. We will likely begin by mapping out our plan discussing portions of the NOAO report entitled Maximizing Science in the Era of LSST. Ben Williams will send around suggested topics and an agenda prior to the telecon.


ACTION ITEMS:

open from previous meetings:

  • Nancy will work with Ben and Russet on a web page that lists the telescope scheduling policies to avoid any future misunderstandings. Status: OPEN. Nancy sent out an email to the institutional schedulers, which then generated some questions that need to be cleared up. A new email and a web page outlining these policies still need to be developed. No change from last month; the scheduling team has been overtaken by the echelle events and hasn't had time to revisit this.
  • Users Committee members should poll their users about interests in AS4 once we distribute the overview document. Status: OPEN. We have not yet gotten the document to distribute.

new from this meeting:

  • There is an ARCSAT usage survey that Nancy sent out to 35m-general and 05m-general mailing lists. Users Committees reps are asked to remind their ARCSAT users to complete this survey (https://goo.gl/forms/4pm2bM1ODUQALEdR2) so that we can gain a better understanding of the role that ARCSAT plays in the research and education programs of ARC member institutions.

Next meeting:

The next Users Committee would normally be on Monday 11 September at 10:30 am MDT. However, this day/time is now bad for at least two Users Committee members, so Nancy will send around a Doodle poll to try and find an optimal meeting day/time.

Note: See TracWiki for help on using the wiki.